# ARECHO: Autoregressive Evaluation via Chain-Based Hypothesis Optimization for Speech Multi-Metric Estimation

### v1

@@ I.2 The Effect of Beam Search
Q-BackgroundNISQA-COLQ-PurposeQ-SpeechImparimentAA-PQQ-ContentRegisterQ-EmotionalVocalizationQ-VoiceTypeQ-PitchRangeQ-EmotionQ-VolumeLevelSSQAQ-GenderSWR/SCRQ-AgeQ-ChannelTypeNISQA-MOSQ-SpeakerCountUTMOSQ-LangAA-CUQ-ClarityNISQA-DISQ-PitchQ-SpeakingStyleLIDQ-EnvQualitySCOREQDNSMOSP.835SE-SI-SNRSingMOSQ-VocComplexityNISQA-NOISE-SARAA-PCUTMOSv2Q-SpeechRateRealLanguageSE-SDRSRMRDNSMOSP.808SE-CI-SDRPAMPLCMOSAA-CENISQA-LOUDSpoofSNISQAReal MOSD-DistanceSCOREQ w. Ref.D-BERTRIR Room SizeSPK-SIMD-BLEUASR-MismatchVoiceMOSReal MOSSARReference Text LengthF0CorrEMO-SIMNoresqaMCDF0RMSEWSSNOMADSTOISDRPredicted Text LengthCI-SDRPESQWERCDSI-SNRRT60CovlCsigNCMCSII-HIGHFWSEGSNRURGENT MOSCSII-LOWLLRVISQOLCSII-MIDCbakSNR SimulationCER
@@
Appendix K Dependency Analysis - Static Dependency Analysis
@@
Inference Enhanced Beam-1 20.58 2.09 1.32 0.84 0.42 0.35 0.72 0.52 0.54 0.51
order-mr 21.66 2.19 1.34 0.83 0.42 0.35 0.72 0.54 0.55 0.53
order-c2f 20.40 2.10 1.31 0.84 0.42 0.35 0.72 0.52 0.54 0.52
Corrupted Beam-1 44.22 2.37 1.29 0.82 0.42 0.35 0.72 0.56 0.56 0.55
order-mr 39.19 2.25 1.23 0.83 0.42 0.35 0.72 0.55 0.57 0.55
order-c2f 42.83 2.32 1.26 0.82 0.42 0.35 0.72 0.56 0.56 0.55
Synthesized Beam-1 4.99 1.13 0.58 0.91 0.45 0.39 0.79 0.67 0.66 0.65
order-mr 4.80 1.12 0.57 0.91 0.45 0.39 0.79 0.67 0.67 0.65
order-c2f 4.65 1.11 0.56 0.91 0.45 0.39 0.79 0.67 0.66 0.65
Fine-tune Enhanced order-mr 22.95 2.25 1.41 0.83 0.42 0.34 0.72 0.54 0.56 0.53
order-c2f 25.40 2.41 1.44 0.82 0.42 0.34 0.72 0.53 0.55 0.53
Corrupted order-mr 39.79 2.30 1.29 0.82 0.42 0.35 0.73 0.56 0.57 0.55
order-c2f 52.72 2.49 1.36 0.82 0.42 0.35 0.72 0.55 0.56 0.54
Synthesized order-mr 4.82 1.13 0.60 0.90 0.45 0.38 0.78 0.65 0.65 0.63
order-c2f 5.36 1.19 0.65 0.90 0.45 0.38 0.77 0.63 0.62 0.60
Appendix L Efficiency Analysis
UniVERSAUniVERSA-TARECHO002,00020002{,}0002 , $0004{,}$ $0006{,}$ $0008{,}$ 0007,668.497668.497{,}668.497 , 668.497,923.347923.347{,}923.347 , 923.346,632.096632.096{,}632.096 , 632.09Avg. Training Time (s / epoch)
@@
Appendix N Broder Impact
@@
Appendix O Acknowledgment

### v2

@@ I.4 Analysis on MOS Prediction
Q-BackgroundNISQA-COLQ-PurposeQ-SpeechImparimentAA-PQQ-ContentRegisterQ-EmoVocalizationQ-VoiceTypeQ-PitchRangeQ-EmotionQ-VolumeLevelSSQAQ-GenderSWR/SCRQ-AgeQ-ChannelTypeNISQA-MOSQ-SpeakerCountUTMOSQ-LangAA-CUQ-ClarityNISQA-DISQ-PitchQ-SpeakingStyleLIDQ-EnvQualitySCOREQDNSMOSP.835SE-SI-SNRSingMOSQ-VocComplexityNISQA-NOISE-SARAA-PCUTMOSv2Q-SpeechRateRealLanguageSE-SDRSRMRDNSMOSP.808SE-CI-SDRPAMPLCMOSAA-CENISQA-LOUDSpoofSNISQAReal MOSD-DistanceSCOREQ w. Ref.D-BERTRIR Room SizeSPK-SIMD-BLEUASR-MismatchVoiceMOSReal MOSSARReference Text LengthF0CorrEMO-SIMNoresqaMCDF0RMSEWSSNOMADSTOISDRPredicted Text LengthCI-SDRPESQWERCDSI-SNRRT60CovlCsigNCMCSII-HIGHFWSEGSNRURGENT MOSCSII-LOWLLRVISQOLCSII-MIDCbakSNR SimulationCER
@@
Appendix L Dependency Analysis - Static Dependency Analysis
@@
Inference Enhanced Beam-1 20.58 2.09 1.32 0.84 0.84 0.70 0.72 0.52 0.54 0.51
order-mr 21.66 2.19 1.34 0.83 0.84 0.70 0.72 0.54 0.55 0.53
order-c2f 20.40 2.10 1.31 0.84 0.84 0.70 0.72 0.52 0.54 0.52
Corrupted Beam-1 44.22 2.37 1.29 0.82 0.84 0.70 0.72 0.56 0.56 0.55
order-mr 39.19 2.25 1.23 0.83 0.84 0.70 0.72 0.55 0.57 0.55
order-c2f 42.83 2.32 1.26 0.82 0.84 0.70 0.72 0.56 0.56 0.55
Synthesized Beam-1 4.99 1.13 0.58 0.91 0.91 0.78 0.79 0.67 0.66 0.65
order-mr 4.80 1.12 0.57 0.91 0.91 0.78 0.79 0.67 0.67 0.65
order-c2f 4.65 1.11 0.56 0.91 0.91 0.78 0.79 0.67 0.66 0.65
Fine-tune Enhanced order-mr 22.95 2.25 1.41 0.83 0.84 0.67 0.72 0.54 0.56 0.53
order-c2f 25.40 2.41 1.44 0.82 0.84 0.67 0.72 0.53 0.55 0.53
Corrupted order-mr 39.79 2.30 1.29 0.82 0.84 0.70 0.73 0.56 0.57 0.55
order-c2f 52.72 2.49 1.36 0.82 0.84 0.70 0.72 0.55 0.56 0.54
Synthesized order-mr 4.82 1.13 0.60 0.90 0.90 0.76 0.78 0.65 0.65 0.63
order-c2f 5.36 1.19 0.65 0.90 0.90 0.76 0.77 0.63 0.62 0.60
Appendix M Efficiency Analysis
UniVERSAUniVERSA-TARECHO02,0002{,}0004,0004{,}0006,0006{,}0008,0008{,}0007,668.497{,}668.497,923.347{,}923.346,632.096{,}632.09Avg. Training Time (s / epoch)
@@
While ARECHO reduces reliance on expensive estimators, it does not replace human listening in high-stakes settings. Automated scores can be misapplied if used out of context (e.g., clinical decisions). We therefore recommend human-in-the-loop verification, transparent reporting of uncertainty, and restricted deployment. Finally, several targets are ordinal; integrating ordinal-aware objectives (e.g., Cao et al. (2020); Gutiérrez et al. (2015); Baccianella et al. (2009)) remains promising future work.
@@
Appendix O Broder Impact
@@
Appendix P ARECHO’s Logo Design
Figure 5: The ARECHO logo: a visual representation of chain-based, autoregressive speech evaluation.
The ARECHO logo was designed to reflect the system’s core principles, autoregessive dependency modeling, multi-metric speech evaluation, and structured reasoning. It consists of two key elements:
  • Waveform Bars: A stylized set of vertical lines representing an audio waveform, symbolizing the raw speech signal input.
  • Interlocked Chain Link: Depicts ARECHO’s classifier chain architecture. It highlights the autoregressive inference procedure and the inter-metric dependencies leveraged during prediction.
These elements are enclosed in a minimal circular boundary to suggest a cohesive and holistic system. Alongside, the bold, uppercase logotype “ARECHO” uses a geometric sans-serif font to convey precision, clarity, and technical strength.
Two visual variants of the logo were created:
  • A light-on-dark version for slides and visual presentations.
  • A symbol-only version for compact branding use (e.g., repository icons or badges).
The design encapsulates ARECHO’s mission: to unify diverse speech evaluation metrics under a structured, dependency-aware modeling framework.
Appendix Q Acknowledgment